OpenStudy (-welp-):

3(x+3)^3/4=81 How do I solve this?

OpenStudy (sshayer):

\[2\left( x+3 \right)^{\frac{ 3 }{ 4 }}=81=3^4\]

OpenStudy (sweetburger):

Divide both sides by 3 then raise both equations to the 4/3 power

OpenStudy (sshayer):

\[\left\{ 2\left( x+3 \right)^{\frac{ 3 }{ 4 }} \right\}^4=\left\{ \left( 3 \right)^4 \right\}^4\]

OpenStudy (sshayer):

or \[\left( x+3 \right)^{\frac{ 3 }{ 4 }}=\frac{ 81 }{ 2 }\] \[\left\{ \left( x+3 \right)^{\frac{ 3 }{ 4 }} \right\}^4=\left( \frac{ 81 }{ 2 } \right)^4\]

OpenStudy (mariamikayla):

3(x+3)^3/4=81 (x+3)^3/4=81/3 (x+3)^3/4=27 (x+3)^3* 1/4= 27 [(x+3)^3]^1/4=27 radical order 4 from (x+3)^3=27 |^4 (x+3)^3= (3^3)^4 (x+3)^3=(3^4)^3 x+3=3^4 x+3=81 x=78

OpenStudy (sshayer):

i wrote2 which is wrong actually it is 3

OpenStudy (sshayer):

\[\left( x+3 \right)^{\frac{ 3 }{ 4 } \times 4}=27^4=\left( 3^3 \right)^4=3^{3 \times 4}=\left( 3^4 \right)^3\] \[x+3=3^4=81,x=81-3=78\]